# Pauli Hamiltonians with an Aharonov-Bohm flux

### v3

@@ 2.2. Self-adjoint extensions
The Friedrichs Hamiltonian $H_{\mathrm{P}}^{(\mathrm{F})}$ is formally recovered taking $\beta=\mbox{``}\infty\mbox{''}\mathds{1}$, i.e., setting all the charges $\mathbf{q}$ equal to zero. Another notable extension is the Krein’s Hamiltonian, i.e., the smallest positive one in form sense. By directed inspection of (2.11), noting that $L(\lambda)$ is positive (see (2.12)), it can be checked that the latter coincides with $H_{\mathrm{P}}^{(0)}$, namely, the extension with $\beta=0$. The same conclusion can be inferred by considering the asymptotics (2.15) and identifying the Krein’s Hamiltonian as the unique extension, besides the Friedrichs’ one, which is homogeneous of degree $-2$ under the $L^{2}$-unitary scaling
@@ 2.4. Symmetries and connection with the Dirac operator
We have already observed that the symmetric operators $H_{\mathrm{P}}$ and $H_{\mathrm{D}}$ defined on $C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{0\};\mathbb{C}^{2})$, are such that $H_{\mathrm{D}}^{2}=H_{\mathrm{P}}$. Now we investigate whether the same property holds for the the respective self-adjoint extensions. More precisely, we aim at determining which extension $H_{\mathrm{P}}^{(\beta)}$ is the square of an extension $H_{\mathrm{D}}^{(\gamma)}$, if any. It turns out that the answer to this question is actually trivial, since the square of all self-adjoint realizations of the Dirac operator coincides with the Fridriechs extension of $H_{\mathrm{P}}$. As a byproduct of the proof, we are also able to classify all the extensions $H_{\mathrm{P}}^{(\beta)}$ such that $\mathscr{D}(H_{\mathrm{P}}^{(\beta)})\subseteq\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$ for some $\gamma\in[0,2\pi)$, i.e., both the first and second order operators are simultaneously well-posed, although in general the latter is not the square of the former.
@@
For any $\gamma\in[0,2\pi)$, there holds
$\left(H_{\mathrm{D}}^{(\gamma)}\right)^{2}=H_{\mathrm{P}}^{(\mathrm{F})}\,.$ (2.70)
Furthermore, $\mathscr{D}(H_{\mathrm{P}}^{(\beta)})\subseteq\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$, for some $\gamma\in[0,2\pi)$, if and only if, for all $\lambda>0$ large enough,
$\begin{cases}\beta^{\ell,0}_{s,\uparrow}=\beta^{\ell,-1}_{s,\downarrow}=\infty\,,\\ [(L(\lambda)+\beta)^{-1}]^{-1,\ell}_{\uparrow,s}=i\cot(\gamma/2)[(L(\lambda)+\beta)^{-1}]^{0,\ell}_{\downarrow,s}\,,\end{cases}\quad\forall\,\ell\in\{-1,0\}\,,\;s\in\{\uparrow,\downarrow\}\,.$ (2.71)
@@
Remark 2.28 (Resonances for Dirac and Pauli operators).
The above result is consistent with the identity $\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}^{2}=H^{(\mathrm{F})}_{\mathrm{P}}$, stated in Proposition 2.26, and with the fact that $H^{(\mathrm{F})}_{\mathrm{P}}$ has no zero-energy resonance (see Proposition 2.21). Indeed, for any given $\gamma\in[0,2\pi)$, even if the spinor $\bm{\psi}_{0}$ found in Proposition 2.27 does satisfy the distributional equation $H_{\mathrm{P}}\bm{\psi}_{0}=(H_{\mathrm{D}}^{\gamma})^{2}\bm{\psi}_{0}=0$, it is not a resonance for the Friedrichs extension $H^{(\mathrm{F})}_{\mathrm{P}}$. This is due to the singular behavior at the origin, making it impossible to satisfy the boundary conditions encoded in $\mathscr{D}(H_{\mathrm{P}}^{(\mathrm{F})})$ (see item $ii)$ in Proposition 2.1 and Definition 2.1).
@@ 3.5. Comparison with the Dirac operator
Consider any of the self-adjoint extensions $H_{\mathrm{P}}^{(\beta)}$, and take $\bm{\psi}\in\mathscr{D}(H_{\mathrm{P}}^{(\beta)})$ as in (2.13). We first prove that if we additionally require that $\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)})^{2}\big{)}$, for some $\gamma\in[0,2\pi)$, then we must have $\bm{\psi}\in\mathscr{D}(H_{\mathrm{P}}^{(\mathrm{F})})$. In other words, our aim is to prove that
$\mathscr{D}\big{(}H_{\mathrm{P}}^{(\beta)}\big{)}\cap\mathscr{D}\big{(}(H_{\mathrm{D}}^{(\gamma)})^{2}\big{)}\subseteq\mathscr{D}\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\big{)}\,.$
By inspection of the proof below, we actually prove a stronger result (see (3.29) and (3.32)), that is,
$\mathscr{D}\big{(}H_{\mathrm{P}}^{(\mathrm{F})}\big{)}\subseteq\mathscr{D}\big{(}(H_{\mathrm{D}}^{(\gamma)})^{2}\big{)}\,,\quad\mbox{for any $\gamma\in[0,2\pi)$.}$ (3.27)
To this aim, we need to show that $\mathbf{q}=\mathbf{0}$, referring to the decomposition in (2.13). This will be achieved imposing
$\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}\,,\qquad H_{\mathrm{D}}^{(\gamma)}\bm{\psi}\in\mathscr{D}\big{(}H_{\mathrm{D}}^{(\gamma)}\big{)}\,.$ (3.28)
We fix the first condition exploiting the characterization of the domain through boundary conditions at the origin, as in (2.69). Recalling (2.8) and (2.9), we observe that, since $\bm{\phi}_{\lambda}\in\mathscr{D}\big{[}Q^{(\mathrm{F})}_{\mathrm{P}}\big{]}$, the spinor vanishes at the origin, so that, for $s\in\{\uparrow,\downarrow\}$,
$c^{s}_{-\alpha}(\bm{\phi}_{\lambda})=0\,,\qquad c^{s}_{\alpha-1}(\bm{\phi}_{\lambda})=0\,.$ (3.29)
Then, thanks to the asymptotics (3.2), it is not hard to see that
$0=c^{\uparrow}_{-\alpha}(\bm{\psi})=2^{-(1-\alpha)}\,\Gamma(\alpha)\,q^{(0)}_{\uparrow}\quad\implies\quad q^{(0)}_{\uparrow}=0\,.$
Similarly, there holds
$0=c^{\downarrow}_{\alpha-1}(\bm{\psi})=2^{-\alpha}\,\Gamma(1-\alpha)\,q_{\downarrow}^{(-1)}\quad\implies\quad q^{(-1)}_{\downarrow}=0\,.$
Moreover, imposing
$c^{\uparrow}_{\alpha-1}(\bm{\psi})=i\cot(\gamma/2)\,\tfrac{2^{1-2\alpha}\,\Gamma(1-\alpha)}{\Gamma(\alpha)}\,c^{\downarrow}_{-\alpha}(\bm{\psi})\,,$
and using again (3.2) we get
$q^{(-1)}_{\uparrow}=i\cot(\gamma/2)\,q^{(0)}_{\downarrow}\,.$ (3.30)
We now need to impose the second condition in (3.28), namely $H_{\mathrm{D}}^{(\gamma)}\bm{\psi}\in\mathscr{D}(H_{\mathrm{D}}^{(\gamma)})$. We preliminarily observe that $\phi_{\lambda}$ can be rewritten as
$\bm{\phi}_{\lambda}=R^{(\mathrm{F})}(-\lambda^{2})\bm{f}\,,\qquad\mbox{for some $\bm{f}=\begin{pmatrix}f^{\uparrow}\\ f^{\downarrow}\end{pmatrix}\in L^{2}(\mathbb{R}^{2};\mathbb{C}^{2})$}\,,$ (3.31)
where $R^{(\mathrm{F})}(-\lambda^{2})$ is the resolvent of the Friedrichs extension $H_{\mathrm{P}}^{(\mathrm{F})}$. Building on this fact and exploiting (2.22), we want to prove that we also have
$c^{s}_{-\alpha}\left(H_{\mathrm{D}}\bm{\phi}_{\lambda}\right)=0\,,\qquad c^{s}_{\alpha-1}\left(H_{\mathrm{D}}\bm{\phi}_{\lambda}\right)=0\,,$ (3.32)
and thus we only need to impose boundary conditions on the singular part of the wave function, in order to find further restrictions on the charges $q^{(-1)}_{\uparrow},q^{(0)}_{\downarrow}$. Recall that, in polar coordinates, $H_{\mathrm{D}}$ reads
$H_{\mathrm{D}}=\begin{pmatrix}0&e^{-i\vartheta}(-i\partial_{r}-\frac{\partial_{\vartheta}+i\alpha}{r})\\ e^{i\vartheta}(-i\partial_{r}+\frac{\partial_{\vartheta}+i\alpha}{r})&0\end{pmatrix}\,.$
Moreover, by (2.22) and (3.31), we get
$(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\uparrow}=\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}-i\tfrac{(k+\alpha)}{r}\right)K_{|k+\alpha|}(\lambda r)\right]\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{|k+\alpha|}(\lambda r^{\prime})f_{\downarrow}^{k}(r^{\prime})\>\tfrac{e^{i(k-1)\vartheta}}{\sqrt{2\pi}}\\ +\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}-i\tfrac{(k+\alpha)}{r}\right)I_{|k+\alpha|}(\lambda r)\right]\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{|k+\alpha|}(\lambda r^{\prime})f_{\downarrow}^{k}(r^{\prime})\>\tfrac{e^{i(k-1)\vartheta}}{\sqrt{2\pi}}\,,$ (3.33)
and
$(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\downarrow}=\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}+i\tfrac{(k+\alpha)}{r}\right)K_{|k+\alpha|}(\lambda r)\right]\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{|k+\alpha|}(\lambda r^{\prime})f_{\uparrow}^{k}(r^{\prime})\>\tfrac{e^{i(k+1)\vartheta}}{\sqrt{2\pi}}\\ +\sum_{k\in\mathbb{Z}}\left[\left(-i\partial_{r}+i\tfrac{(k+\alpha)}{r}\right)I_{|k+\alpha|}(\lambda r)\right]\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{|k+\alpha|}(\lambda r^{\prime})f_{\uparrow}^{k}(r^{\prime})\>\tfrac{e^{i(k+1)\vartheta}}{\sqrt{2\pi}}\,.$ (3.34)
Notice that in the above formula boundary terms coming from the $r$-dependent extremes of integration cancel out. Recall that the verification of the boundary conditions require the evaluation of the linear functionals (2.65). Let us prove that
$c^{\uparrow}_{-\alpha}\big{(}H_{\mathrm{D}}\bm{\phi}_{\lambda}\big{)}=0\,,$ (3.35)
the other conditions in (3.32) being treated similarly. Taking the angular average and using basic properties of Bessel functions, we find
$\left\langle(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\uparrow}\right\rangle(r)=i\lambda\left(K_{\alpha}(r)\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{1+\alpha}(\lambda r^{\prime})f_{\downarrow}^{1}(r^{\prime})-I_{\alpha}(r)\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{1+\alpha}(\lambda r^{\prime})f_{\downarrow}^{1}(r^{\prime})\right).$ (3.36)
We are now led to analyze the asymptotic behavior of the last term on the r.h.s. of the above formula, as $r\to 0^{+}$. To this aim, recall the asymptotics of Bessel functions $I_{\nu}(t)\sim t^{\nu}$ and $K_{\nu}(t)\sim t^{-\nu}$ as $t\to 0^{+}$. Then, there holds
$\left|\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}I_{1+\alpha}(\lambda r^{\prime})f_{\downarrow}^{1}(r^{\prime})\right|\leqslant\left\|f_{\downarrow}\right\|_{2}\left(\int^{r}_{0}\,\mathrm{d}r^{\prime}r^{\prime}(I_{1+\alpha}(\lambda r^{\prime}))^{2}\right)^{1/2}\leqslant Cr^{2+\alpha}\,,$
where $C>0$ is a constant independent of $r$. Moreover, we get
$\left|\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}K_{1+\alpha}(\lambda r^{\prime})f_{\downarrow}^{1}(r^{\prime})\right|\leqslant\left\|f_{\downarrow}\right\|_{2}\left(\int^{+\infty}_{r}\,\mathrm{d}r^{\prime}r^{\prime}(K_{1+\alpha}(\lambda r^{\prime}))^{2}\right)^{1/2}\leqslant Cr^{-\alpha}\,.$
Combining the above observations, we find
$\left\langle(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\uparrow}\right\rangle(r)=O(1)\,,\qquad\mbox{as $r\to 0^{+}$,}$
and then
$c^{\uparrow}_{-\alpha}\big{(}H_{\mathrm{D}}\bm{\phi}_{\lambda}\big{)}=\lim_{r\to 0^{+}}r^{\alpha}\left\langle(H_{\mathrm{D}}\bm{\phi}_{\lambda})_{\uparrow}\right\rangle(r)=0\,.$
As already remarked, the other conditions in (3.32) follow by similar arguments.
Now, observe that
$H_{\mathrm{D}}\bm{\psi}=H_{\mathrm{D}}\bm{\phi}_{\lambda}+q^{(-1)}_{\uparrow}H_{\mathrm{D}}\bm{G}^{(-1)}_{\lambda,\uparrow}+q^{(0)}_{\downarrow}H_{\mathrm{D}}\bm{G}^{(0)}_{\lambda,\downarrow}=H_{\mathrm{D}}\bm{\phi}_{\lambda}+i\lambda\left(q^{(-1)}_{\uparrow}\bm{G}^{(0)}_{\lambda,\downarrow}+q^{(0)}_{\downarrow}\bm{G}^{(-1)}_{\lambda,\uparrow}\right)\,,$ (3.37)
as we have

### v4

@@ 2.2. Self-adjoint extensions
The Friedrichs Hamiltonian $H_{\mathrm{P}}^{(\mathrm{F})}$ is formally recovered taking $\beta=\mbox{``}\infty\mbox{''}\mathds{1}$, i.e., setting all the charges $\mathbf{q}$ equal to zero. Another distinguished extension is the Krein’s Hamiltonian, i.e., the smallest positive one in form sense. By directed inspection of (2.11), noting that $L(\lambda)$ is positive (see (2.12)), it can be inferred that the latter coincides with $H_{\mathrm{P}}^{(0)}$, namely, the extension with $\beta=0$. It is also noteworthy that the Friedrichs’ and the Krein’s Hamiltonians are homogeneous of degree $-2$ under the $L^{2}$-unitary scaling
@@
The same property is shared by the other Hamiltonians $H_{\mathrm{P}}^{(\beta)}$ which coincide with either the Friedrichs or the Krein one in each of the channels $(s,\ell)$, $s\in\{\uparrow,\downarrow\}$ and $\ell\in\{0,-1\}$, separately.
@@ 2.4. Symmetries and connection with the Dirac operator
We have already observed that the symmetric operators $H_{\mathrm{P}}$ and $H_{\mathrm{D}}$ defined on $C^{\infty}_{c}(\mathbb{R}^{2}\setminus\{0\};\mathbb{C}^{2})$, are such that $H_{\mathrm{D}}^{2}=H_{\mathrm{P}}$. Now we investigate whether the same property holds for the the respective self-adjoint extensions. More precisely, we aim at determining which extension $H_{\mathrm{P}}^{(\beta)}$ is the square of an extension $H_{\mathrm{D}}^{(\gamma)}$, if any. It turns out that only for specific values of the parameter $\gamma$ and for specific choices of $\beta=\big{(}\beta_{ss^{\prime}}^{(\ell\ell^{\prime})}\big{)}\in\mathrm{M}_{4,\mathrm{Herm}}(\mathbb{C})$ there actually holds $(H_{\mathrm{D}}^{\gamma})^{2}=H^{(\beta)}_{\mathrm{P}}$.
@@
The identity
$(H_{\mathrm{D}}^{\gamma})^{2}=H^{(\beta)}_{\mathrm{P}}$ (2.70)
holds if and only if one of the following two alternatives is realized:
  1. (a)
    $\gamma=0$, $\beta_{\uparrow\uparrow}^{(00)}=\beta_{\downarrow\downarrow}^{(00)}=\beta_{\downarrow\downarrow}^{(-1-1)}=\infty$ and $\beta_{\uparrow\uparrow}^{(-1-1)}=0$ ($\beta_{ss^{\prime}}^{(\ell\ell^{\prime})}$ arbitrarily chosen otherwise), i.e., $H^{(\beta)}_{\mathrm{P}}$ coincides with the Krein extension in the $(\uparrow,-1)$ channel and with the Friedrichs extension in all the other channels;
  2. (b)
    $\gamma=\pi$, $\beta_{\uparrow\uparrow}^{(00)}=\beta_{\uparrow\uparrow}^{(-1-1)}=\beta_{\downarrow\downarrow}^{(-1-1)}=\infty$ and $\beta_{\downarrow\downarrow}^{(00)}=0$ ($\beta_{ss^{\prime}}^{(\ell\ell^{\prime})}$ arbitrarily chosen otherwise), i.e., $H^{(\beta)}_{\mathrm{P}}$ coincides with the Krein extension in the $(\downarrow,0)$ channel and with the Friedrichs extension in all the other channels.
Remark 2.27.
The Dirac extensions corresponding to $\gamma=0$ and $\gamma=\pi$ are two distinguished ones. They are indeed the only scale covariant realizations, i.e., homogeneous of degree $-1$ under scaling, for generic values of $\alpha\in(0,1)$. [^1] Namely, for all $\alpha\neq 1/2$. The case $\alpha=1/2$ is indeed exceptional, since in this instance any Dirac Hamiltonian $H_{\mathrm{D}}^{(\gamma)}$, $\gamma\in[0,2\pi)$, exhibits scale covariance. This facts has implications at the dynamical level. Indeed, in [CDYZ] the authors remark that taking
$\begin{cases}\gamma=\pi\,,\qquad\mbox{for }\alpha\in(0,1/2]\,,\\ \gamma=0\,,\qquad\mbox{for }\alpha\in(1/2,1)\,,\end{cases}$ (2.71)
one gets the only self-adjoint realizations $H_{\mathrm{D}}^{(\gamma)}$ whose domain is contained in $H^{1/2}(\mathbb{R}^{2}\,;\mathbb{C}^{2})$, a fact which allows to identify them as distinguished ones. Notice that here we are referring to our parametrization, which is slightly different from the one adopted in [CDYZ]. Some dispersive estimates are then proven in [CDYZ], for the choice of parameters (2.71). It is also mentioned therein that for other extensions such estimates seem to fail.
@@
Remark 2.29.
The above result is coherent with Proposition 2.21, concerning the resonances for Pauli AB Hamiltonians $H^{(\beta)}_{\mathrm{P}}$(see also Remark 2.22). By direct inspection, it is readily seen that the latter coincide with those of $H^{\gamma}_{\mathrm{D}}$ in (2.72), for the distinguished choices of the parameters $\gamma\in[0,2\pi)$ and $\beta\in\mathrm{M}_{4,\mathrm{Herm}}(\mathbb{C})$ described in the preceding Proposition 2.26, so that $(H^{\gamma}_{\mathrm{D}})^{2}=H^{(\beta)}_{\mathrm{P}}$.
@@ 3.5. Comparison with the Dirac operator
The proof of the thesis will be achieved in various steps. Without loss of generality, in what follows we fix $\lambda=1$, omitting the dependence on the spectral parameter for simplicity of notation. We will repeatedly use the expansion as $r\to 0^{+}$ given in (3.2) (see [GR07, Eq. 6.521.3]), alongside with the basic identities